# Extreme Forward Scattering Observed in Disk-Averaged Near-Infrared Phase Curves of Titan

### v1

@@ 4.2 Notable Spectral Features
Notable features in apparent emission are seen in Figure 2 at 3.31 m and 4.54 m. For context, Figure 10 shows two high-quality examples of whole disk images of Titan at the wavelengths with these notable emission features. The 3.31 m emission primarily comes from the dayside limb while the 4.54 m is structured and concentrated near the southern pole. Figure 11 highlights the spectral region encompassing these features from the heatmap in Figure 4, demonstrating how the features maintain a near-constant with increasing phase angle. García-Comas et al. (2011) explain that the 3.31 m feature is non-local thermodynamic equilibrium (non-LTE) emission from upper-atmospheric methane driven by absorption of solar radiation. While both CO and CH3D contribute to non-LTE emission near 4.54 m, where the fundamental of CH3D gives rise to a sharp feature at 4.54 m (Baines et al., 2006), emission from these species is not strong enough to explain a sharp increase in seen in some cubes at this wavelength. Thus we suspect that a subset of cubes are affected by residual instrument or calibration issues at this wavelength, such as an order-sorting filter change, hot pixel, or other impediment in channel 317.
@@ 4.3 Phase Curve Variability
The 2 m phase curve in Figure 5 reveals a spread whose strength increases with decreasing phase angle. This behavior is apparent in phase curves from all wavelengths with deep atmosphere and/or surface sensitivity, thus indicating that processes in the deep atmosphere or on the surface are driving the phase curve variability. For example, the visibility of Xanadu — a large, spectrally distinct region on the surface of Titan centered at about 100S longitude and 10S latitude (Coustenis et al., 1995; Lellouch et al., 2004; Negrão et al., 2006) — should introduce variability at surface-sensitive wavelengths.
An indication that weather plays some role in causing variability in phase curves with deep atmospheric sensitivity is highlighted by flybys 253TI-255TI, flyby 264TI, and flyby 273TI (hereafter referred to as groups 1, 2, and 3, respectively), which dominate the brightest, low-phase observations in the 2 m curve in Figure 5. Looking at visualizations of cubes from each group reveals a set of clouds above Titan’s north pole, surrounded by a ring of circumpolar clouds at about 50N (13). All three groups took place during early- to mid-spring 2017. The cloud ring is most easily viewed in early group 3 cubes, while the polar cloud can easily be seen in cubes from groups 1 and 2. The location and shapes of cloud structures are in agreement with the findings of Yahn et al. (2025), who discovered that, from September 2016 to September 2017, clouds on Titan were concentrated around 0-120W longitude, 50N latitude, and had large aspect ratios indicating long, thin structures.
4.4 Application to Exoplanetary Atmospheres
The presence of aerosols in Titan’s atmosphere evidently dominates near-infrared reflected-light phase curves of the moon. It is reasonable, then, to expect similar mechanisms to occur in the atmospheres of hazy exoplanets. Models (Hu et al., 2013; Adams et al., 2019; Gao et al., 2020) and laboratory results (He et al., 2018; Hörst et al., 2018) have shown that exoplanets with a wide range of parameters are capable of hosting aerosol-forming atmospheres, though currently the strongest detections of aerosols come from warm/hot giant planets (e.g. Estrela et al., 2021; Malsky et al., 2025). Observations of Titan can serve as a rare opportunity to study the phase curves of cool, terrestrial exoplanets with thick atmospheres, and especially the impact of atmospheric aerosols thereon.
@@ 5 Conclusions
We developed a pipeline that takes Cassini images of Titan and automatically determines the disk-averaged of Titan at all VIMS-IR bands. Our dataset of 4,604 images spanned a phase angle range of 2 to 165. We produced phase curves of Titan in the near-infrared regime (0.88–5.12 m) that show strong non-Lambertian effects.
Our near-infrared reflectance spectrum of Titan at 2 is a useful approximation to Titan’s geometric albedo. This observation extends important previous results at primarily optical wavelengths (Karkoschka, 1998) and also agrees with very recent, high-quality near-full phase spectral observations at near-infrared and infrared wavelengths from James Webb Space Telescope (Nixon et al., 2025).
Phase curves of Titan at wavelengths with surface sensitivity show enhanced backscattering at low phase angles as well as increased variability at these phase angles due to clouds in the lower/deep atmosphere. Additionally, spectra show evidence of non-LTE emission, particularly at 3.31 m (CH4) even when averaged over the entire disk.
@@ 6 Acknowledgments
The present work was initially supported by a seed grant from the Arizona Astrobiology Center, led by CC. All authors acknowledge support through an award from NASA’s Exoplanets Research Program (No. 80NSSC25K7149). CC and TR conceived of this study, which was further developed by CC, TR, JB, and LM. All pipeline materials, results, and figures were created by CC. CC wrote all early drafts of this manuscript, with later editing support from TR, JB, and LM. We thank A. García Muñoz for sharing optical phase curve data for Titan.

### v3

@@ 4.2 Notable Spectral Features
Notable features in apparent emission are seen in Figure 2 at 3.31 m and 4.54 m. For context, Figure 9 shows two high-quality examples of whole disk images of Titan at the wavelengths with these notable emission features. The 3.31 m emission primarily comes from the dayside limb while the 4.54 m emission is structured and concentrated near the southern pole. Figure 10 highlights the spectral region encompassing these features from the color-contour diagram in Figure 3, demonstrating how both features maintain a near-constant with increasing phase angle, which further suggests this feature is due to emission rather than phase-dependent forward scattering. Garc´ıa-Comas et al. (2011) explain that the 3.31 m feature is non-local thermodynamic equilibrium (non-LTE) emission from upper-atmospheric methane driven by absorption of solar radiation. The isotropic emission from this non-LTE source results in a phase curve shape that is distinctly less-structured than at wavelengths dominated by absorption or scattering processes, as seen in Figure 3. This emission is prevalent enough to produce a spectral feature even when averaged over the disk of Titan. While both CO and CH3D contribute to non-LTE emission near 4.54 m, where the fundamental of CH3D gives rise to a sharp feature at 4.54 m (Baines et al., 2006), emission from these species is not strong enough to explain a sharp increase in seen in some cubes at this wavelength. Thus we suspect that a subset of cubes are affected by residual instrument or calibration issues at this wavelength, such as an order-sorting filter change, hot pixel, or other impediment in channel 317.
@@ 4.3 Phase Curve Variability
The 2 m phase curve in Figure 4 reveals a spread whose strength increases with decreasing phase angle. The spread far exceeds the anticipated error in disk-averaged values, especially at these longer wavelengths that do not suffer from saturation issues. This behavior is apparent in phase curves from all wavelengths with deep atmosphere and/or surface sensitivity, thus indicating that processes in the deep atmosphere or on the surface are driving the phase curve variability. For example, the visibility of Xanadu — a large, spectrally distinct region on the surface of Titan centered at about 100S longitude and 10S latitude (Coustenis et al., 1995; Lellouch et al., 2004; Negrão et al., 2006) — should introduce variability at surface-sensitive wavelengths. However, we found no substantive correlation between the sub-spacecraft coordinates of Cassini at the time of cube acquisitions and the relative of the cubes (Figure 12).
An indication that weather plays some role in causing variability in phase curves with deep atmospheric sensitivity is highlighted by flybys 253TI-255TI, flyby 264TI, and flyby 273TI (hereafter referred to as groups 1, 2, and 3, respectively), which dominate the brightest, low-phase observations in the 2 m curve in Figure 4. Looking at visualizations of cubes from each group reveals a set of clouds above Titan’s north pole, surrounded by a ring of circumpolar clouds at about 50N (Figure 11). All three groups took place during early- to mid-spring 2017. The cloud ring is most easily viewed in early group 3 cubes, while the polar cloud can easily be seen in cubes from groups 1 and 2. The location and shapes of cloud structures are in agreement with the findings of Yahn et al. (2025), who discovered that, from September 2016 to September 2017, clouds on Titan were concentrated around 0-120W longitude, 50N latitude, and had large aspect ratios indicating long, thin structures.
4.4 Applications and Future Work
The shape of Titan’s phase curves are directly related to the physical properties of its hazes, such as their size, opacities, and single scattering albedo (Tomasko et al., 2008; Doose et al., 2016; Garc´ıa Muñoz et al., 2017). Their wavelength-properties have been deduced previously from similar measurements at visible and some near-infrared wavelengths. The findings of this work could be used to extend the results of past studies by further constraining haze parameters at near-infrared wavelengths. This, in turn, would find applications in determining Titan’s energy budget, modeling its atmospheric thermodynamics, or validating models thereof (Garc´ıa Muñoz et al., 2017).
The presence of aerosols in Titan’s atmosphere evidently dominates near-infrared reflected-light phase curves. It is reasonable, then, to expect similar mechanisms to occur in the atmospheres of hazy exoplanets. Models (Hu et al., 2013; Adams et al., 2019; Gao et al., 2020) and laboratory results (He et al., 2018; Hörst et al., 2018) have shown that exoplanets with a wide range of atmospheric conditions are capable of hosting aerosols, and many strong detections of aerosols come from warm/hot giant planets (e.g. Estrela et al., 2021; Malsky et al., 2025). Observations of Titan can serve as a rare opportunity to study the phase curves of cool, terrestrial exoplanets with thick atmospheres, and especially the impact of atmospheric aerosols thereon.
@@ 5 Conclusions
We developed a pipeline that takes Cassini images of Titan and automatically determines the disk-averaged of Titan at all VIMS-IR bands. Our dataset of 4,492 images spanned a phase angle range of 2 to 165. We produced phase curves of Titan in the near-infrared regime (0.88–5.12 m) that show strong non-Lambertian effects.
Our near-infrared reflectance spectrum of Titan at 2 is a useful approximation to Titan’s geometric albedo. This observation extends important previous results at primarily optical wavelengths (Karkoschka, 1998) and also agrees with very recent, high-quality near-full phase spectral observations at near-infrared and infrared wavelengths from James Webb Space Telescope (Nixon et al., 2025, their Figure 4b).
Phase curves of Titan at wavelengths with surface sensitivity show enhanced backscattering at low phase angles as well as increased variability at these phase angles due to clouds in the lower/deep atmosphere. Additionally, spectra show evidence of a feature at 3.31 m attributable to CH4 non-LTE emission. While this feature has been observed and studied before in spatially-resolved observations (Baines et al., 2005; Garc´ıa-Comas et al., 2011), we found that this feature is observable even when averaged over the entire disk.
@@ 6 Acknowledgments
The present work was initially supported by a seed grant from the Arizona Astrobiology Center, led by CC. All authors acknowledge support through an award from NASA’s Exoplanets Research Program (No. 80NSSC25K7149). CC and TR conceived of this study, which was further developed by CC, TR, JB, and LM. All pipeline materials, results, and figures were created by CC. CC wrote all early drafts of this manuscript, with later editing support from TR, JB, and LM. We thank A. García Muñoz for sharing optical phase curve data for Titan. We also thank the two anonymous reviewers for their insightful comments and suggestions for improvement.
7 Appendix - Data Acquisition and Calibration
Data used in this work were acquired using the PDS Image Atlas. The data were part of the ”VIMS Observations from the Cassini Tour of the Saturn System” dataset (Brown & VIMS Science Team, 2020). Data were selected from the dataset by requiring that they target Titan. The ISIS3 package is developed by the United States Geological Survey and can be accessed from their GitHub page.